OpenStudy (anonymous):

HELP PLEASE ASAP??? a 1,600.00 principle earns 7% interest, compound semiannually twice per year. After 33 years, what is the balance in the account?

OpenStudy (anonymous):

i think it is 3696

OpenStudy (anonymous):

A=1600(1+0.07/2) ^66 A=1600(1.035)^66 A=*1600*9.684=15,494.69 round up to 15,494.70 I found it already but thanks anyways! :)